Find the general solution of the differential equation xy = y + (x2 + y2 y(4) = 3
Answer #1

Given differential equen is xy = y + J x² + y2 y= dy dr » y= y + x² + y2 Leta y=va then, y= utav then (1) = v+xv wat Jx²+=) =) 2 log I vt Jl+v2 | = loglal + loghitel ; logi sel = coabitrary constant 1v+Jit 11+52 1 -\1121 # (v+ $1+vz)- x2 (squarin